# An RSK correspondence for cylindric tableaux

### 10 Mar 2026 v1

@@ Abstract.
This paper establishes an analogue of the Robinson–Schensted correspondence for cylindric tableaux. In particular, for any pair of positive integers $(d,L)$, we construct a bijection between permutations that avoid the patterns $d\cdots 1(d+1)$ and $1\cdots(L+1)$ and pairs of $(d,L)$-cylindric standard Young tableaux with a common shape. This arises as a special case of a Knuth-type generalization involving cylindric semistandard tableaux and a further generalization involving oscillating tableaux. Using these results, we construct several other bijections and derive enumerative consequences involving cylindric tableaux and pattern-avoiding permutations. For example, we give asymptotics for the number of permutations in $S_{n}$ that avoid the patterns $d\cdots 1(d+1)$ and $1\cdots(L+1)$ as $n\to\infty$.
@@ 1.1. Overview
Although the connection between cylindric tableaux and pattern-avoiding permutations that we establish in this paper is new, there are some natural predecessors of our work in the literature. In particular, Bloom and Saracino [4] defined a certain map using growth diagrams which is essentially the $d$-RS correspondence. Their work does not involve cylindric tableaux, however. On the other hand, Elizalde [10] used growth diagrams to reformulate a correspondence of Neyman [26] which involves cylindric tableaux but not permutations. Our work shows that these growth diagrams are closely related. In Section 11 we will give a direct comparison of the local rules used in our growth diagrams and those used by Elizalde and Bloom and Saracino. We will also generalize the results of Elizalde and Neyman in Section 7.
@@ 1.2. Outline
The remainder of this paper is organized as follows. In Section 2 we establish our notation and the necessary background on partitions, Young diagrams, and pattern avoidance, and we formally define various notions of cylindric tableaux. In Section 3 we define growth diagrams and state our main results. In Section 4 we derive some immediate corollaries of our main results (in particular, the cylindric RS and cylindric RSK correspondences). In Sections 5 and 6 we give the proofs of the main results. In Section 7 we show how our growth diagrams can be modified to give a different correspondence involving skew cylindric tableaux. This generalizes work of Elizalde and Neyman.
The remaining sections give various auxiliary consequences of the correspondences described in Sections 3–7. In Section 8, we discuss cylindric conjugation and cylindric row-strict tableaux, and how conjugation can be combined with the correspondences. In Section 9, we derive some additional bijections involving fillings of Young diagrams. In Section 10, we give several results on permutations in $S_{n}$ that avoid the patterns $d\cdots 1(d+1)$ and $1\cdots(L+1)$ including an asymptotic formula for the number of such permutations as $n\to\infty$. In Section 11, we discuss how to local rules used in our growth diagrams are related to local rules used by Elizalde [10], and Bloom and Saracino [4]. In Section 12, we describe continuous piecewise-linear generalizations of our correspondences.
@@ 2.3. Semistandard Young tableaux
It is well-known (see [23, p. 5]) that any semistandard Young tableau $T$ corresponds to a filling of the Young diagram of $\mathrm{sh}(T)$ with positive entries that are weakly increasing along rows from left to right and strictly increasing along columns from bottom to top. To obtain this filling, the entry $i$ is placed in all the cells that are in the Young diagram of $\lambda^{(i)}$ but not in the Young diagram of $\lambda^{(i-1)}$. Hence, the number of cells with entry $i$ is equal to $\mathrm{wt}(T)_{i}$. An example is shown in Figure 4(a).
@@ 2.4. Oscillating tableaux
We call a semistandard $w$-oscillating tableau a $d$-semistandard $w$-oscillating tableau if all of its constituent partitions are $d$-partitions. Moreover, we call it a $(d,L)$-cylindric semistandard $w$-oscillating tableau if we can replace all the interlacing relations in the definition with $(d,L)$-interlacing relations. Given a $d$-semistandard oscillating tableau $T=(\lambda^{(0)},\ldots,\lambda^{(r)})$, we define its minimum cylindric width, denoted $\mathrm{MCW}_{d}(T)$, to be the smallest $L$ such that $T$ is $(d,L)$-cylindric. Equivalently, $\mathrm{MCW}_{d}(T)=\max_{1\leq i\leq r}\mathrm{MCW}_{d}(\lambda^{(i-1)},\lambda^{(i)})$.
We define the analogue of weight in the oscillating setting as follows. Given a semistandard $w$-oscillating tableau $T=(\emptyset=\lambda^{(0)},\lambda^{(1)},\ldots,\lambda^{(r)}=\emptyset)$, we define $\mathrm{wt}^{+}(T)$ to be the vector of nonnegative integers whose $i$th entry is $\mathrm{wt}^{+}(T)_{i}=|\lambda^{(a_{i})}|-|\lambda^{(a_{i}-1)}|$ where $a_{i}$ is the index of the $i$th occurrence of ‘++’ in $w$. We define $\mathrm{wt}^{-}(T)$ to be the vector of nonnegative integers whose $i$th entry is $\mathrm{wt}^{-}(T)_{i}=|\lambda^{(b_{i}-1)}|-|\lambda^{(b_{i})}|$ where $b_{i}$ is the index of the $i$th-to-last occurrence of ‘-’ in $w$.
@@ 2.5. Standard tableaux
The various kinds of tableaux that we defined above have the word “semistandard” in their names. We replace “semistandard” with “standard” in the special case where the associated weight vectors consist of all $1$’s. For example, a standard Young tableau is a semistandard Young tableau $T$ such that $\mathrm{wt}(T)=(1,1,\ldots,1)$. A standard $w$-oscillating tableau is a semistandard $w$-oscillating tableau $T$ such that $\mathrm{wt}^{+}(T)=(1,1,\ldots,1)$ and $\mathrm{wt}^{-}(T)=(1,1,\ldots,1)$.
@@ 3.1. Growth diagrams and local rules
RSK local rule: It is required that $\mu\succ\kappa\prec\nu$ and $\mu\succ\rho\prec\nu$. Moreover,
@@
$d$-RSK local rule: It is required that $\mu\succ\kappa\prec\nu$ and $\mu\succ\rho\prec\nu$. Also, all four partitions are $d$-partitions, and it is required that $m=0$ or $\kappa_{d}=0$. Moreover,
@@
The RSK local rule is well known (cf. [20, Sec. 4.1], [35, Sec. 3.1]). Growth diagrams satisfying this local rule encode the RSK correspondence in a natural way as we will recall shortly. The $d$-RSK local rule is new, but from the definition one can see that it is a “cyclic” modification of the RSK local rule. We will show that $d$-RSK growth diagrams encode a bijective correspondence that we call the $d$-RSK correspondence. An example of a $d$-RSK growth diagram is shown in Figure 6.
@@ 3.2. Growth diagram correspondences
Theorem 3.2 (Oscillating RSK correspondence; see [20, Thm. 7]).
@@ 3.3. Properties of the oscillating RSK and $d$-RSK correspondences
The fact that these two reflection operations correspond to each other under the oscillating RSK or $d$-RSK correspondence is equivalent to the statement that the correspondences commute with reflection. Properly interpreting this statement requires viewing the correspondences as unified bijections over all choices of Young diagrams.
Alternatively, if $F$ is a Young diagram that is symmetric about $y=x$, then this property also holds at the level of a single Young diagram. In this special case, a consequence is that symmetric fillings of $F$ correspond to “palindromic” tableaux (i.e. those that are invariant under reversal).
The first property stated in Proposition 3.4 is more straightforward. We will refer to this property as the weight-preserving property of RSK and $d$-RSK. Although we did not specifically define the notion of weight for fillings, it should be clear why we have chosen this name.
@@ 3.4. Chains in growth diagrams
This theorem is a growth diagram analogue of a classical theorem of Schensted [32, Thm. 1]. It has a generalization due to Greene [14] (see [20, Thm. 8]) that gives a complete description of the partitions in the growth diagram in terms of chains, but we will not need this generalization.
@@ 4. Cylindric correspondences
Recently, Huh, Kim, Krattenthaler, and Okada [15] discovered other combinatorial expressions for the number of $(d,L)$-cylindric standard Young tableaux of a given size. It would be interesting to establish a direct bijection between their expressions and the one given above. The main results of Huh et al. are certain symmetric function identities, which are affine analogues of bounded Littlewood identities (see [15, Sec. 1]). Even in the non-affine setting these bounded Littlewood identities are not yet well understood from a bijective perspective (see [11]).
@@ 7. Skew $d$-RSK correspondence
In this section we construct another correspondence that we call the oscillating skew $d$-RSK correspondence. This is closely related to the oscillating $d$-RSK correspondence although neither is a generalization of the other. At the end of the section we will explain how this correspondence generalizes prior work of Neyman [26]; Elizalde [10]; and Courtiel, Elvey Price, and Marcovici [8].
@@ 7.2. Skew $d$-RSK growth diagrams
Skew $d$-RSK local rule: Let $\kappa,\mu,\nu,\rho$ denote the $d$-staircases assigned to the corners of the cell (as labelled in Figure 5), and let $m$ denote the entry. It is required that $\mu\succ\kappa\prec\nu$ and $\mu\succ\rho\prec\nu$ and that $m=0$. Moreover,
@@ 7.4. Special cases and connections to prior work
The oscillating skew $d$-RSK correspondence can be restricted in various ways to obtain other correspondences. In particular, by placing an upper bound $L$ on the minimum cylindric width of the tableaux, we obtain a correspondence involving $(d,L)$-cylindric tableaux. Moreover, if we take the type sequences $w$ and $v$ to be of the form $w=+^{r}-^{c}$ and $v=-^{c}+^{r}$, we obtain a non-oscillating version of this cylindric correspondence between pairs of $(d,L)$-cylindric semistandard Young tableaux. Such a correspondence was constructed previously by Neyman [26] using a variant of Schensted insertion. Our work gives an oscillating generalization of Neyman’s correspondence.
In the special case of standard tableaux, Neyman’s bijection was reformulated by Elizalde [10, Sec. 5] using growth diagrams. Elizalde’s local rule coincides with ours in this case. We discuss this further in Section 11. Elizalde also used the growth diagram construction to describe an oscillating correspondence just as we have done (see [10, Sec. 6]). He showed that this oscillating correspondence is equivalent to a certain correspondence constructed by Courtiel, Elvey Price, and Marcovici [8] involving walks in a $d$-dimensional simplex. Hence, our work generalizes these correspondences to the semistandard setting.
It is worth noting that neither Elizalde nor Neyman considered $(d,L)$-cylindric correspondences as restrictions of a broader correspondence that is independent of $L$ as we have done. This is a relatively minor distinction, but we believe that the perspective we have taken here clarifies the nature of these correspondences and their relationships to one another.
@@
The skew $d$-RSK correspondence can also be used to give bijective proofs of certain identities involving skew Schur functions and cylindric Schur functions (see [26, Thm. 5.14, Cor. 5.21] for the cylindric case). In the non-cylindric setting, the resulting skew Schur function identity is equivalent to a certain relation involving sums of Littlewood–Richardson coefficients. This relation was first proved by Coquereaux and Zuber [7, Thm. 1] using representation theory.
@@ 8.2. Cylindric analogues
There is a subtle distinction between the interlacing and cointerlacing which is worth emphasizing. For cointerlacing, if $\lambda$ and $\mu$ are $d$-staircases
@@
$\lambda\prec_{(d,L)}\mu\implies\Big(\lambda\text{ and }\mu\text{ are }(d,L)\text{-staircases, and }\lambda\prec\mu\Big).$
@@
Properties (i)–(iii) of this proposition be found in work of Goodman and Wenzl [13, Lem. 1.5], although their definitions are somewhat different than ours. Their version is also only stated for $(d,L)$-partitions. Property (iv) of this proposition is important because it gives the desired duality between interlacing and cointerlacing for $(d,L)$-staircases.
@@ 8.3. Duality and RSK-type correspondences
To state this result we must define oscillating tableaux in the row-strict setting. It suffices to say that we can take all of the definitions from the semistandard oscillating setting and carry them over to the row-strict oscillating setting by replacing interlacing with cointerlacing and $(d,L)$-interlacing with $(d,L)$-cointerlacing. We can now state the following result.
@@
Taking $L\to\infty$ in Theorem 8.9 leads to nontrivial enumerative consequences involving skew $d$-row-strict $w$-oscillating tableaux. For example, the number of such tableaux of a given shape and weight is independent of the choice of type sequence $w$ (assuming we fix the number of occurrences of ‘++’ and ‘-’). Hence, we obtain a statement about non-cylindric objects via cylindric correspondences.

### 13 Mar 2026 v2

@@ Abstract.
This paper establishes an analogue of the Robinson–Schensted correspondence for cylindric tableaux. In particular, for any pair of positive integers $(d,L)$, we construct a bijection between permutations that avoid the patterns $d\cdots 1(d+1)$ and $1\cdots(L+1)$ and pairs of $(d,L)$-cylindric standard Young tableaux with a common shape. This arises as a special case of a Knuth-type generalization involving cylindric semistandard tableaux and a further generalization involving oscillating tableaux. Using these results, we construct several other bijections and derive enumerative consequences involving cylindric tableaux and pattern-avoiding permutations. For example, we give an asymptotic for the number of permutations in $S_{n}$ that avoid the patterns $d\cdots 1(d+1)$ and $1\cdots(L+1)$ as $n\to\infty$.
@@ 1.1. Overview
Although the connection between cylindric tableaux and pattern-avoiding permutations that we establish in this paper is new, there are some natural predecessors of our work in the literature. In particular, Bloom and Saracino [4] defined a certain map using growth diagrams which is essentially the $d$-RS correspondence. Their work does not involve cylindric tableaux, however. On the other hand, Elizalde [10] used growth diagrams to reformulate a correspondence of Neyman [27] which involves cylindric tableaux but not permutations. Our work shows that these growth diagrams are closely related. In Section 11 we will give a direct comparison of the local rules used in our growth diagrams and those used by Elizalde and Bloom and Saracino. In Section 7 we will generalize Neyman’s correspondence.
@@ 1.2. Outline
The remainder of this paper is organized as follows. In Section 2 we establish our notation and the necessary background on partitions, Young diagrams, and pattern avoidance, and we formally define various notions of cylindric tableaux. In Section 3 we define growth diagrams and state our main results. In Section 4 we derive some immediate corollaries of our main results (in particular, the cylindric RS and cylindric RSK correspondences). In Sections 5 and 6 we give the proofs of the main results. In Section 7 we show how our growth diagrams can be modified to construct a different correspondence involving skew cylindric tableaux. This is a generalization of a correspondence of Neyman [27].
The remaining sections give various auxiliary consequences of the correspondences described in Sections 3–7. In Section 8, we discuss cylindric conjugation and cylindric row-strict tableaux, and how conjugation can be combined with the correspondences. In Section 9, we derive some additional bijections involving fillings of Young diagrams. In Section 10, we prove several results on permutations in $S_{n}$ that avoid the patterns $d\cdots 1(d+1)$ and $1\cdots(L+1)$ such as an asymptotic formula for the number of such permutations as $n\to\infty$. In Section 11, we discuss how the local rules used in our growth diagrams are related to local rules used by Bloom and Saracino [4], and Elizalde [10]. In Section 12, we describe continuous piecewise-linear generalizations of our correspondences.
@@ 2.3. Semistandard Young tableaux
It is well known (see [24, p. 5]) that any semistandard Young tableau $T$ corresponds to a filling of the Young diagram of $\mathrm{sh}(T)$ with positive entries that are weakly increasing along rows from left to right and strictly increasing along columns from bottom to top. To obtain this filling, the entry $i$ is placed in all the cells that are in the Young diagram of $\lambda^{(i)}$ but not in the Young diagram of $\lambda^{(i-1)}$. Hence, the number of cells with entry $i$ is equal to $\mathrm{wt}(T)_{i}$. An example is shown in Figure 4(a).
@@ 2.4. Oscillating tableaux
A semistandard $w$-oscillating tableau is called a $d$-semistandard $w$-oscillating tableau if all of its constituent partitions are $d$-partitions. Moreover, we call it a $(d,L)$-cylindric semistandard $w$-oscillating tableau if we can replace all the interlacing relations in the definition with $(d,L)$-interlacing relations. Given a $d$-semistandard oscillating tableau $T=(\lambda^{(0)},\ldots,\lambda^{(r)})$, we define its minimum cylindric width, denoted $\mathrm{MCW}_{d}(T)$, to be the smallest $L$ such that $T$ is $(d,L)$-cylindric. Equivalently, $\mathrm{MCW}_{d}(T)=\max_{1\leq i\leq r}\mathrm{MCW}_{d}(\lambda^{(i-1)},\lambda^{(i)})$.
The analogue of weight in the oscillating setting is defined as follows. Given a semistandard $w$-oscillating tableau $T=(\emptyset=\lambda^{(0)},\lambda^{(1)},\ldots,\lambda^{(r)}=\emptyset)$, we associate two vectors of nonnegative integers, $\mathrm{wt}^{+}(T)$ and $\mathrm{wt}^{-}(T)$, to it. The $i$th entry of $\mathrm{wt}^{+}(T)$ is $\mathrm{wt}^{+}(T)_{i}=|\lambda^{(a_{i})}|-|\lambda^{(a_{i}-1)}|$ where $a_{i}$ is the index of the $i$th occurrence of ‘++’ in $w$. The $i$th entry of $\mathrm{wt}^{-}(T)$ is $\mathrm{wt}^{-}(T)_{i}=|\lambda^{(b_{i}-1)}|-|\lambda^{(b_{i})}|$ where $b_{i}$ is the index of the $i$th-to-last occurrence of ‘-’ in $w$. (Note the difference in the ordering between the two definitions.)
@@ 2.5. Standard tableaux
For each of the different kinds of tableaux defined above, we replace the word “semistandard” with “standard” in the special case where the associated weight vectors consist of all $1$’s. For example, a standard Young tableau is a semistandard Young tableau $T$ such that $\mathrm{wt}(T)=(1,1,\ldots,1)$. A standard $w$-oscillating tableau is a semistandard $w$-oscillating tableau $T$ such that $\mathrm{wt}^{+}(T)=(1,1,\ldots,1)$ and $\mathrm{wt}^{-}(T)=(1,1,\ldots,1)$.
@@ 3.1. Growth diagrams and local rules
RSK local rule: It is required that $\mu\succ\kappa\prec\nu$ and $\mu\prec\rho\succ\nu$. Moreover,
@@
$d$-RSK local rule: It is required that $\mu\succ\kappa\prec\nu$ and $\mu\prec\rho\succ\nu$. Also, all four partitions are $d$-partitions, and it is required that $m=0$ or $\kappa_{d}=0$. Moreover,
@@
The RSK local rule is well known (cf. [21, Sec. 4.1], [36, Sec. 3.1]). Growth diagrams satisfying this local rule encode the RSK correspondence in a natural way as we will recall shortly. The $d$-RSK local rule is new, but from the definition one can see that it is a “cyclic” modification of the RSK local rule. We will show that $d$-RSK growth diagrams encode a bijective correspondence that we call the $d$-RSK correspondence. An example of a $d$-RSK growth diagram is shown in Figure 6.
@@ 3.2. Growth diagram correspondences
Theorem 3.2 (Oscillating RSK correspondence; see [21, Thm. 7]).
@@ 3.3. Properties of the oscillating RSK and $d$-RSK correspondences
To refer to the fact that these two reflection operations correspond to each other under the oscillating RSK or $d$-RSK correspondence, we say that the correspondences commute with reflection. Properly interpreting this statement requires viewing the correspondences as unified bijections over all choices of Young diagrams.
Alternatively, this commutativity property is also meaningful at the level of a single Young diagram $F$ if $F$ happens to be symmetric about $y=x$. In this special case, this property has an interesting consequence: symmetric fillings of $F$ correspond to “palindromic” tableaux (i.e. those that are invariant under reversal).
The first property stated in Proposition 3.4 is more straightforward. We will refer to this as the weight-preserving property of RSK and $d$-RSK. Although we did not specifically define the notion of weight for fillings, it should be clear why we have chosen this name.
@@ 3.4. Chains in growth diagrams
This theorem is a growth diagram analogue of a classical theorem of Schensted [33, Thm. 1]. It has a generalization due to Greene [15] (see [21, Thm. 8]) that gives a complete description of the partitions in the growth diagram in terms of chains, but we will not need this generalization.
@@ 4. Cylindric correspondences
Recently, Huh, Kim, Krattenthaler, and Okada [16] discovered other combinatorial expressions for the number of $(d,L)$-cylindric standard Young tableaux of a given size. It would be interesting to establish a direct bijection between their expressions and the one given above. The main results of Huh et al. are certain symmetric function identities, which are affine analogues of bounded Littlewood identities (see [16, Sec. 1]). Even in the non-affine setting these bounded Littlewood identities are not yet well understood from a bijective perspective (see [12]).
@@ 7. Skew $d$-RSK correspondence
In this section we construct another correspondence that we call the oscillating skew $d$-RSK correspondence. This is closely related to the oscillating $d$-RSK correspondence although neither is a generalization of the other. At the end of the section we will explain how this correspondence relates to prior work of Neyman [27]; Elizalde [10]; and Courtiel, Elvey Price, and Marcovici [8].
@@ 7.2. Skew $d$-RSK growth diagrams
Skew $d$-RSK local rule: Let $\kappa,\mu,\nu,\rho$ denote the $d$-staircases assigned to the corners of the cell (as labelled in Figure 5), and let $m$ denote the entry. It is required that $\mu\succ\kappa\prec\nu$ and $\mu\prec\rho\succ\nu$ and that $m=0$. Moreover,
@@ 7.4. Special cases and connections to prior work
The oscillating skew $d$-RSK correspondence can be restricted in various ways to obtain other correspondences. In particular, by placing an upper bound $L$ on the minimum cylindric width of the tableaux, we obtain a correspondence involving $(d,L)$-cylindric tableaux. Moreover, if we take the type sequences $w$ and $v$ to be of the form $w=+^{r}-^{c}$ and $v=-^{c}+^{r}$, we obtain a non-oscillating version of this cylindric correspondence between pairs of skew $(d,L)$-cylindric semistandard Young tableaux. Such a correspondence was constructed previously by Neyman [27] using a variant of Schensted insertion. Our work gives an oscillating generalization of Neyman’s correspondence.
In the case of cylindric standard tableaux, Neyman’s bijection was reformulated by Elizalde in [10] using growth diagrams. Our local rule agrees with Elizalde’s in this setting (see Section 11). Elizalde also used the growth diagram construction to describe an oscillating cylindric correspondence. He showed that this correspondence is equivalent to a certain correspondence constructed by Courtiel, Elvey Price, and Marcovici [8] involving walks in a $d$-dimensional simplex.
We may summarize the contributions of this section as follows. Firstly, we showed that a simple modification of the $d$-RSK local rule leads to skew correspondences in a natural way. Secondly, the cylindric correspondences that were considered by Neyman and Elizalde are actually restrictions of broader correspondences that are independent of the $L$ parameter. Thirdly, the cylindric correspondences that were considered by Neyman and Elizalde can be naturally extended to the semistandard oscillating setting.
@@
After the first version of this paper was posted to the arXiv, Elizalde informed us that he had independently extended his work in [10] to the semistandard setting. This will appear in [11]. The local rule he uses for this is identical to our skew $d$-RSK local rule apart from the difference in perspective mentioned above (i.e. our correspondences are independent of the $L$ parameter).
Remark 7.4.
The skew $d$-RSK correspondence can also be used to give bijective proofs of certain identities involving skew Schur functions and cylindric Schur functions (see [27, Thm. 5.14, Cor. 5.21] for the cylindric case). In the non-cylindric setting, the resulting skew Schur function identity is equivalent to a certain relation involving sums of Littlewood–Richardson coefficients. This relation was first proved by Coquereaux and Zuber [7, Thm. 1] using representation theory.
@@ 8.2. Cylindric analogues
There is a subtle point that is worth emphasizing regarding a certain distinction between interlacing and cointerlacing. For cointerlacing, if $\lambda$ and $\mu$ are $d$-staircases
@@
$\lambda\prec_{(d,L)}\mu\implies\Big(\lambda\text{ and }\mu\text{ are }(d,L)\text{-staircases, and }\lambda\prec\mu\Big),$
@@
Properties (i)–(iii) of this proposition be found in work of Goodman and Wenzl [14, Lem. 1.5], although their definitions are somewhat different than ours. Their version is also only stated for $(d,L)$-partitions. Property (iv) of this proposition is important because it gives the desired duality between interlacing and cointerlacing for $(d,L)$-staircases.
@@ 8.3. Duality and RSK-type correspondences
To state this result we must define oscillating tableaux in the row-strict setting. It suffices to say that we can take all of the definitions from the semistandard oscillating setting and carry them over to the row-strict oscillating setting by replacing interlacing with cointerlacing and $(d,L)$-interlacing with $(d,L)$-cointerlacing. With these definitions in place, we can state the following result.
@@
It should also be noted that taking $L\to\infty$ in Theorem 8.9 leads to a nontrivial enumerative consequence involving skew $d$-row-strict $w$-oscillating tableaux. That is, the number of such tableaux of a given shape and weight is independent of the choice of type sequence $w$ (assuming we fix the number of occurrences of ‘++’ and ‘-’). This latter statement does not involve cylindric objects, but we have obtained it as a consequence of cylindric correspondences.